If 5L of a gas is at 1.08 atm . Supposing the volume was 10L what will the volume be at a standard temperature


1
Expert's answer
2021-02-07T19:07:23-0500

"p_1V_1=p_2V_2,"

"p_2=p_1\frac{V_1}{V_2},"

"p_2=1.08\cdot \frac{5}{10}=0.54~atm."